OpenStudy (anonymous):

6/x-9e^x find the antiderivative of f(x)

OpenStudy (tkhunny):

\(\dfrac{6}{x} - 9e^{x}\;\;?\) Which part troubles you, the logarithm or the exponential?

OpenStudy (anonymous):

the 9e^x. I'm quite confident for the first part the answer would come out as 6x^-1? right?

OpenStudy (tkhunny):

?? Isn't it already \(6x^{-1}\)? Hint 1 \(\dfrac{d}{dx}e^{x} = e^{x}\) Hint 2 \(\dfrac{d}{dx}\ln(x) = 1/x\) <== Careful with this one, it doesn't quite work the other direction.

OpenStudy (anonymous):

so it'd be 9(1/x)? I'm looking through my notebook right now and I see a similar problem to it but I don't have a problem that has an e^x only just a e

OpenStudy (abb0t):

\(\int\limits e^{-x}dx = -e^{-x}+C\), you can check this by taking the derivative of \(e^{-x}\)

OpenStudy (abb0t):

\(-e^{-x}\)

OpenStudy (anonymous):

but there's the 9 in front? So I don't think I can do anything like that to the e

OpenStudy (abb0t):

9 is a constant and you can simply factor it out.

OpenStudy (abb0t):

You can also check this by taking the derivaitve of \(9[-e^{-x}]\) also.

OpenStudy (anonymous):

isnt the derivative of e^x just 1 though?

OpenStudy (tkhunny):

So, figure out that \(\int e^x\;dx = e^{x} + C\). Let the chain rule and the negative go!

OpenStudy (abb0t):

No. Use chain rule. derivative is simply \(e^x\)

OpenStudy (anonymous):

so the derivative of 9 is 9x and e^x is just e^x (1) but the 1 doesn't really matter so i dont put that so its 9xe^x?

OpenStudy (abb0t):

No! \(\frac{ d }{ dx }(-9e^{-x})\) = \(9[-1 \times e^{-x} \times \frac{ d }{ dx }(-x)] = 9[(-1 \times -1) \times e^{-x}] = 9e^{-x}\), which as you can see is just the anti-derivative.

OpenStudy (anonymous):

so 9e^-x is the anti-derivative of the 2nd part o.O?

OpenStudy (abb0t):

No, \(-9e^{-x}\) is the anti-derivative of \(9e^{-x}\)

OpenStudy (anonymous):

ok